OpenStudy (anonymous):

Two charged balloons 2.1*10^-1 meters apart have a repulsive force of 1.3*10^-1 newtons. The charge on one balloon is 8.2*10^-7 coulombs. What is the charge on the other balloon?

OpenStudy (anonymous):

\[F=k\frac{ Q_1Q_2}{ r^2 }\]

OpenStudy (anonymous):

F=Force between the charges k=Coulomb's Law constant (9.0 x 10^9) Q1=quantity of charge on object 1. Q2=quantity of charge on object 2. r=distance of separation between the two objects

OpenStudy (anonymous):

NB: the value of k is dependent upon the medium the objects are immersed in......in the case of air, the value of the constant is what i have given to you.....

OpenStudy (anonymous):

\[Q_2=\frac{ Fr^2 }{ kQ_1 }\] \[Q_2=\frac{ 1.3 \times 10^{-1}\times (2.1 \times10^{-1})^2 }{ 9.0 \times 10^{9} \times 8.2\times 10^{-7} }\] \[Q_2=7.8 \times 10^{-7}\]